OpenStudy (anonymous):

Two objects (49.0 and 18.0 kg) are connected by a massless string that passes over a massless, frictionless pulley. The pulley hangs from the ceiling. Find (a) the acceleration of the objects and (b) the tension in the string.

OpenStudy (anonymous):

Equation For Object1 { M1= 18kg} T-m1g= m1a => T = m1a + m1g .........Eq1 Equation For Object2 { M2= 49 Kg} T- m2g = -m2a => T = m2g - m2a........ Eq 2 Equating eq 1 and 2... we get m1a + m1g = m2g - m2a =>m1a +m2a = m2g - m1g => 18a + 49a = 49g - 18g => 67a = 31g => a = 31g/67 So acceleration is 31g/67 And Tension = 18 * 31g/67 +18a